OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a18 = 97, a20 = 281

OpenStudy (anonymous):

arithmetic sequence meaning that its increasing by a certain number so first determine what number its increasing by a20-a18/(20-18) then determine what a0 is then write in terms of an=a0 + n (increasing number)

OpenStudy (anonymous):

1st, insert the value like just now u'll get 2 equations like these :- 97 = a + 17d 281 = a + 19d then use simultenous equation :)

OpenStudy (anonymous):

an = -1467 - 92(n + 1) an = -1467 + 92(n + 1) an = -1467 + 92(n - 1) an = -1467 + 92(n - 2) These are my choices. I put the third one, but I'm not sure if it's right or not. I learned all of this last year, but I completely forgot my lessons.

OpenStudy (anonymous):

u can find d 1st by eliminating a, then find a, after obtaining both value, insert it in general formula ... wait, lemme try :)

OpenStudy (anonymous):

yeap, third 1

OpenStudy (anonymous):

Yay :) Thanks, honey!

OpenStudy (anonymous):

no problem :D